OpenStudy (anonymous):

Finding Combinations Evaluate each expression 25C7

OpenStudy (anonymous):

\[_{25}C_7=\frac{\overline{25\times 24\times 23\times 22\times 21\times 20\times 19}^{7 numbers}}{7\times 6\times 5\times 4\times 3\times 2\times 1}\] cancel first, multiply last

OpenStudy (anonymous):

the arithmetic is somewhat tedious, most calculators can do it for you so can the web http://www.wolframalpha.com/input/?i=25+choose+7

OpenStudy (anonymous):

\[_{25} C_7=\frac{ 25! }{7!(25-7)! }\] You can cancel most of the terms if you want to do it by hand.

OpenStudy (anonymous):

As a matter of fact, @satellite73 has already canceled most of the factors which can be removed.

OpenStudy (anonymous):

btw why aren't you watching the knicks?

OpenStudy (anonymous):

I think my brother is.